Find the mean and variance of following data 10,60,50,30,40,20


1
Expert's answer
2021-08-31T16:16:58-0400

"Mean = \frac{10+60+50+30+40+20}{6}= 35 \\

Variance = \frac{(10-35)^2+(60-35)^2+(50-35)^2+(30-35)^2+(40-35)^2+(20-35)^2}{6} \\

= \frac{625+625 + 225 + 25 + 25 + 225}{6} = 291.66"